### v2

# ![[Uncaptioned image]](extracted/6365550/Fig/teaser.png) Syzygy of Thoughts: Improving LLM CoT with the Minimal Free Resolution

Abstract.
Chain-of-Thought (CoT) prompting enhances the reasoning of large language models (LLMs) by decomposing problems into sequential steps, mimicking human logic and reducing errors. However, complex tasks with vast solution spaces and vague constraints often exceed the capacity of a single reasoning chain. Inspired by Minimal Free Resolution (MFR) in commutative algebra and algebraic geometry, we propose Syzygy of Thoughts (SoT)—a novel framework that extends CoT by introducing auxiliary, interrelated reasoning paths. SoT captures deeper logical dependencies, enabling more robust and structured problem-solving. MFR decomposes a module into a sequence of free modules with minimal rank, providing a structured analytical approach to complex systems. This method introduces the concepts of ”Module”, ”Betti numbers”,”Freeness”, ”Mapping”, ”Exactness” and ”Minimality”, enabling the systematic decomposition of the original complex problem into logically complete minimal subproblems while preserving key problem features and reducing reasoning length. We tested SoT across diverse datasets (e.g., GSM8K, MATH) and models (e.g., GPT-4o-mini, Qwen2.5), achieving inference accuracy that matches or surpasses mainstream CoTs standards. Additionally, by aligning the sampling process with algebraic constraints, our approach enhances the scalability of inference time in LLMs, ensuring both transparent reasoning and high performance. Our code will be publicly available at https://github.com/dlMARiA/Syzygy-of-thoughts
1. Introduction
The rapid advancements in natural language processing (Vaswani et al., 2017; Devlin et al., 2019; Radford et al., 2018; Brown et al., 2020) have demonstrated outstanding reasoning capabilities across various tasks (Zhao et al., 2023b; Min et al., 2023). In recent years, the Chain of Thought (CoT) prompting method (Wei et al., 2022) has emerged as an effective strategy for simulating human reasoning by generating intermediate reasoning steps, significantly enhancing model performance in solving complex problems (Chen et al., 2025; Qu et al., 2025; Sui et al., 2025; Wei et al., 2025; Xia et al., 2024; Liu et al., 2025). The core advantage of the CoT framework lies in its ability to decompose problems into a series of explicit and easily interpretable intermediate steps, thereby improving the transparency and explainability of the reasoning process. This explicit stepwise reasoning not only facilitates in-depth analysis and debugging of model outputs but also provides a clear logical pathway for solving complex problems.
However, despite its success in many tasks, CoT exhibits inherent limitations when dealing with high-dimensional, nonlinear, and abstract logical problems (Wang et al., 2022a; Shi et al., 2023; Jin et al., 2024). First, traditional CoT methods often rely on fixed or heuristic decomposition strategies, which struggle to capture essential details and intrinsic connections in multidimensional problems (Wang et al., 2022a). Second, due to the lack of systematic structural planning in intermediate steps, the reasoning process tends to include redundant computations and repetitive paths, leading to increased computational resource consumption and inference latency (Jin et al., 2024). Although existing improvements have alleviated these issues to some extent, introducing techniques such as multi-path exploration, voting mechanisms, and hierarchical decomposition strategies to enhance robustness (Yao et al., 2023a; Besta et al., 2024; Wang et al., 2022b; Zhou et al., 2022; Yao et al., 2023b), however, they have yet to fully address logical incompleteness and the inadequate decomposition of high-dimensional information. As a result, when handling highly complex reasoning tasks, model robustness and accuracy still face significant challenges.
To address these challenges, we propose a novel reasoning framework—Syzygy of Thoughts (SoT). This method draws inspiration from the concept of MFR in algebraic geometry and computational algebra (Eisenbud, 2013; Rossi and Sharifan, 2009; Botbol et al., 2021), systematically decomposing and reconstructing problems by constructing minimal free module sequences. The incorporation of MFR theory ensures a more rigorous and efficient problem decomposition process while preserving the core characteristics of the problem, effectively reducing redundant computations and logical inconsistencies. Specifically, SoT retains the stepwise reasoning advantages of CoT while introducing algebraic tools such as ”module”, ”Betti numbers”, ”freeness”, ”mapping”, ”exactness” and ”minimality” to deeply explore and systematically express the intrinsic structure of complex problems. This approach provides a novel theoretical perspective for analyzing high-dimensional and multivariable problems.
To validate the effectiveness of SoT, we conducted extensive experiments across diverse datasets, including GSM8K, MATH, and others, spanning mathematical reasoning and high-dimensional problem domains. These datasets, with their complex structures and multifaceted features, provide a robust testbed for evaluating reasoning frameworks in real-world scenarios. Experimental results show that SoT achieves inference accuracy that matches or surpasses mainstream CoTs standards, demonstrating its superior performance and practical potential.
Finally, the SoT framework exhibits multiple advantages, including an intuitive and transparent reasoning path and improved interpretability of intermediate information. The main contributions of this work are as follows:
Pioneering the integration of MFR into LLM reasoning, enabling structured decomposition of complex problems.
Developing a novel framework that builds on CoT’s strengths while substantially improving reasoning clarity and structural efficiency.
Providing theoretical insights and empirical evidence that SoT significantly enhances performance in high-dimensional reasoning tasks.
2. Related Work
2.1. Advanced Reasoning Techniques in LLMs
The reasoning capabilities of LLMs have advanced significantly with breakthroughs in prompting paradigms, notably CoT prompting proposed by Wei et al.(Wei et al., 2022). CoT decomposes complex problems into stepwise solutions, mimicking human reasoning, yet struggles with multidimensional or nonlinear tasks like abstract logic and advanced mathematics. To address these limitations, variants such as Zero-shot-CoT(Kojima et al., 2022), Self-Consistency CoT (Wang et al., 2022b), and Auto-CoT (Zhang et al., 2022) have emerged, alongside verification-based approaches like VerifyCoT (Zhao et al., 2023a) and CoF-CoT (Nguyen et al., 2023), yielding measurable improvements.
Structured prompting strategies further enhance reasoning. Least-to-Most Prompting (LtM)(Zhou et al., 2022) breaks tasks into manageable subproblems, reducing single-step errors. Programmatic methods, including Program of Thought (PoT)(Chen et al., 2022), Chain of Code (CoC)(Li et al., 2023), and Buffer of Thought (BoT)(Yang et al., 2025), integrate discrete variables and procedural steps. Building on these, Algorithm of Thought (AoT) (Sel et al., 2023) iteratively synthesizes and refines solutions, boosting token efficiency.
Tree- and graph-based paradigms offer alternative exploration. Tree-of-Thought (ToT)(Yao et al., 2023a) employs a hierarchical structure to explore multiple decision paths, though exponential growth limits deep efficiency. Graph-of-Thought (GoT)(Besta et al., 2024) extends this with flexible graph configurations, enabling path recalibration and information aggregation. Meanwhile, Skeleton-of-Thought (SoT)(Ning et al., 2023) generates high-level outlines before detailing, balancing speed and quality, while recursive methods like Self-Refine(Madaan et al., 2023) and Step-Back Prompting (Zheng et al., 2023) refine outputs iteratively for higher accuracy.
2.2. Minimal Free Resolution
MFR is a core tool in homological algebra (Fieldsteel and Nagel, 2021) and algebraic geometry (Botbol et al., 2021), used to analyze the algebraic structure of modules (Eisenbud, 2013; Rossi and Sharifan, 2009), revealing properties such as rank, symmetry, and relationships (Chen, 2011). In computational algebraic geometry, MFR analyzes singularities and invariants of algebraic varieties through syzygies (Evans and Griffith, 1981), optimizing the complexity and efficiency of Gröbner basis computations (Eisenbud et al., 2001; Stewart, 1993; Weispfenning, 1992; Capani et al., 1997).
In Topological Data Analysis (TDA), MFR accelerates persistent homology computations (Wasserman, 2018), enhancing the stability of topological analysis for high-dimensional data. In physics and bioinformatics, MFR is used to analyze Calabi-Yau manifold singularities (Polchinski, 1994), gauge field structures (Hitchin, 2003), and the modular properties of gene regulatory networks (Li et al., 2011).
The advantages of MFR in optimizing computational complexity and revealing algebraic structures provide a new perspective for its integration into the study of CoT in LLM. In CoT, structured reasoning steps are required, but often face challenges of inefficiency and redundancy. We innovatively introduce MFR into CoT by decomposing the symbolic dependencies between reasoning steps, eliminating redundant calculations, optimizing intermediate processes, and enhancing transparency. This approach opens up a more efficient and structured technical pathway for LLM applications in solving complex problems.
3. Method
3.1. Preliminaries
In algebraic geometry and homological algebra, minimal free resolution is an essential tool for studying the structure and properties of modules, especially when dealing with finitely generated modules over a ring.
Key Properties: In algebraic geometry and homological algebra, a ”syzygy” refers to the relations among the relations that generate a module. A syzygy is a relation among the generators. For example, if a module is defined by generators , then a syzygy is a relation of the form:
where . These relations themselves can form a module, which is called the ”first syzygy module.” A module over a ring is an algebraic structure similar to a vector space, but the scalars come from instead of a field. A module is called free if it has a basis, i.e., for some dimension (finite or infinite). Free modules are analogous to free vector spaces in linear algebra. A module is finitely generated if there exists a finite set such that . A MFR of a finitely generated module over is an exact sequence:
Where each is free (), the matrices of the maps contain no units, and the resolution is ”minimal,” meaning the number of free generators is the smallest possible. This resolution provides a standardized way to describe the structure of , reflecting both its complexity and its homological properties. A MFR is characterized by its minimality, where each uses the smallest number of generators, and the matrices in the maps contain no units. It is unique up to isomorphism, meaning any two MFRs of are naturally isomorphic. The -th Betti number of , defined as , counts the number of free generators in and reflects the complexity of the resolution.
3.2. Minimal Free Resolution enhanced CoT
As shown in Figure 3, the SoT reasoning paradigm employs the MFR framework to model the structure of CoT reasoning paths in LLMs. Each reasoning step—such as problem interpretation, task decomposition, subgoal chaining, and final inference—is abstracted as generators or relations within modules, reinterpreting CoT as algebraic objects with generative structures and latent dependencies, rather than merely a flat sequence of symbols.
![[Uncaptioned image]](extracted/6365550/Fig/modul.png)
Module: Initial Complex Problem
Complex problems often exhibit high-dimensional structures, intricate logical dependencies, and ambiguous constraints, rendering direct solutions intractable. In this initial module, we conceptualize the problem as a unified entity of interdependent substructures, necessitating systematic decomposition. The goal is not immediate resolution but to unravel latent complexity, identify core logical components, and establish a foundation for subsequent modular reasoning. This stage ensures that downstream processes operate within well-defined boundaries, setting the stage for structured analysis.
![[Uncaptioned image]](extracted/6365550/Fig/betti.png)
Betti Numbers: Quantifying Complexity
To measure decomposition complexity, we introduce Betti numbers as a quantitative metric, where each number reflects the count of auxiliary conditions at a reasoning level. Higher Betti numbers indicate greater intricacy, signaling potential optimization opportunities. We leverage LLMs to minimize these numbers by regenerating or filtering conditions, streamlining the decomposition process. Betti numbers thus guide reasoning efficiency, balancing structural complexity with computational simplicity.
![[Uncaptioned image]](extracted/6365550/Fig/freeness.png)
Freeness
: Auxiliary Conditions
Freeness involves generating auxiliary conditions to simplify problems and clarify logical relationships. By introducing intermediate variables, hypotheses, or formalized constraints, complex problems are partitioned into smaller, independent subproblems. This reduces dimensionality and provides clear entry points for resolution, enhancing the tractability of subsequent steps while maintaining logical coherence.
![[Uncaptioned image]](extracted/6365550/Fig/mapping.png)
Mapping: Solving Strategies
Effective problem-solving requires systematic strategies that map auxiliary conditions to actionable reasoning paths. These paths must exhibit (i) directness, minimizing redundant operations, and (ii) logical soundness, grounding each step in explicit premises). LLMs iteratively solve subproblems along these paths, linking current conditions to new conclusions, ensuring continuity and interpretability throughout the reasoning chain.
![[Uncaptioned image]](extracted/6365550/Fig/exactness.png)
Exactness: Logical Completeness
Logical completeness is critical for rigor, preventing gaps or leaps in the reasoning chain. Each step must derive from explicit premises, with implicit assumptions formalized to avoid invalid conclusions. LLMs verify consistency across auxiliary conditions, subproblems, and the final solution, systematically reviewing the chain against original constraints. This ensures a coherent and reliable framework, minimizing errors from incomplete logic.
![[Uncaptioned image]](extracted/6365550/Fig/mini.png)
Minimality: Optimal Solutions
Minimality optimizes efficiency by deriving solutions with the fewest auxiliary conditions and simplest strategies, without sacrificing correctness. Redundant conditions are filtered, and reasoning paths are pruned of unnecessary steps. LLMs prioritize straightforward approaches over complex transformations, yielding concise, clear solutions. This reduces overall complexity, enhancing both computational efficiency and solution transparency.
After introducing the above six concepts, LLMs can be directly stimulated in a conversational manner. Figure 4 illustrates the conversational framework of these six concepts. First, complex reasoning problems are defined as a free module. Then, through the Betti number and auxiliary conditions of degrees of freedom, the complex problem is reduced and decomposed. By analyzing both the module and degrees of freedom simultaneously, the model can construct multiple mappings to solve the problem. Finally, minimality and accuracy are used to select the optimal solution.
3.3. Syzygy of Thoughts
The process of LLM reasoning for complex problems can be viewed as solving a complex reasoning problem through MFR. The entire process is formalized via a MFR sequence, where the structured representation of the reasoning process is a set of mappings induced by the LLM, forming the following sequence:
Moreover, we define the intermediate space as a free module constructed from basic reasoning units:
where:
- is the reasoning space (LLM Reasoning Space), - are the basic units (Tokens) used in the reasoning construction, - Each is induced by the LLM, representing a single-step logical deduction in the CoT.
3.3.1. SoT Reasoning.
The Algorithm 1 and Figure 3 explain the entire process of SoT. The SoT framework begins with an initial complex problem , characterized by high-dimensional structure, intricate logical dependencies, and ambiguous constraint relationships. We treat as a unified representation of interdependent substructures, setting the stage for systematic decomposition. To analyze the inherent complexity of , we introduce Betti numbers as quantitative indicators of structural complexity. Specifically, the Betti number corresponds to the number of auxiliary conditions required for decomposition, with higher values suggesting deeper structural intricacy. These values guide the subsequent generation of auxiliary components and indicate potential optimization points.
We then introduce the notion of Freeness, which refers to the generation of auxiliary conditions that serve to simplify and structurally reorganize the problem. These conditions—denoted as —include intermediate hypotheses, latent constraints, or explicit sub-conclusions, and are generated via autoregressive reasoning using LLMs conditioned on previously established states. Each decomposes the original problem into finer-grained logical units, reducing its dimensional complexity and establishing solvable subspaces.
To resolve the decomposed problem, the framework constructs a set of mappings , each corresponding to a distinct reasoning path. These mappings are designed to satisfy two essential properties: directness, ensuring that each inference step avoids redundant operations, and logical soundness, ensuring that every conclusion is firmly grounded in its premises. Each mapping is generated using LLMs to traverse from auxiliary conditions to intermediate or final conclusions, thereby forming a continuous and interpretable reasoning chain.
To ensure logical completeness (Exactness) of the reasoning process, all steps within each reasoning path are verified for inferential closure. Implicit assumptions are formalized, and the consistency of auxiliary conditions, mappings, and final conclusions is evaluated. This validation step leverages the LLM’s reasoning and scoring capabilities via a scoring function . Finally, the optimal reasoning sequence is parsed using structured pattern-matching strategies (e.g., regular expressions, JSON parsing) to extract the final solution.
Method | Math Reasoning | Gene. | Multitask | Temporal | Log. | ||||
GSM8K | SVAMP | MultiArith | ASDiv | AQUA | MMLU | BBH | Date | CLUTRR | |
GPT-4o-mini | |||||||||
CoT | 85.1% | 84.4% | 99.2% | 97.0% | 65.0% | 63.1% | 66.3% | 51.8% | 65.9% |
CoT-SC (n=5) | 90.1% | 86.0% | 99.5% | 98.5% | 70.9% | 67.3% | 69.2% | 54.9% | 72.4% |
SoT (Ours) | 96.0% | 92.2% | 99.7% | 99.8% | 75.6% | 75.2% | 72.8% | 75.2% | 75.7% |
Qwen2.5-Coder-7B-Instruct | |||||||||
CoT | 77.2% | 82.4% | 92.3% | 92.0% | 60.6% | 55.1% | 47.1% | 31.0% | 20.1% |
CoT-SC (n=5) | 80.2% | 84.1% | 95.0% | 95.0% | 62.2% | 56.3% | 49.3% | 32.9% | 21.0% |
SoT (Ours) | 89.1% | 90.6% | 97.0% | 99.8% | 63.3% | 57.1% | 57.3% | 36.2% | 26.3% |
Qwen2.5-VL-72B-Instruct | |||||||||
CoT | 86.1% | 86.9% | 98.8% | 98.0% | 81.1% | 80.1% | 77.3% | 75.2% | 70.1% |
CoT-SC (n=5) | 89.1% | 88.2% | 99.3% | 98.4% | 83.9% | 82.9% | 79.0% | 78.0% | 75.0% |
SoT (Ours) | 96.0% | 95.8% | 99.7% | 99.2% | 89.4% | 84.3% | 85.3% | 80.2% | 78.9% |
Gemma-3-27b-it | |||||||||
CoT | 83.1% | 85.9% | 91.9% | 98.5% | 80.3% | 70.8% | 70.7% | 76.9% | 65.3% |
CoT-SC (n=5) | 87.1% | 87.0% | 92.3% | 99.2% | 85.4% | 73.2% | 73.2% | 80.2% | 66.4% |
SoT (Ours) | 96.0% | 95.8% | 99.7% | 99.2% | 89.4% | 84.3% | 85.3% | 80.2% | 78.9% |
Gemma-3-12b-it | |||||||||
CoT | 83.2% | 79.0% | 90.4% | 97.7% | 68.9% | 68.1% | 64.6% | 77.7% | 49.0% |
CoT-SC (n=5) | 86.1% | 81.0% | 93.3% | 98.0% | 71.7% | 70.6% | 66.7% | 80.2% | 52.2% |
SoT (Ours) | 92.1% | 92.5% | 96.1% | 99.2% | 77.2% | 72.3% | 69.1% | 82.5% | 55.0% |
4. Experiments
To evaluate the performance of SoT, we designed a series of experiments, including applicability evaluation (Sec4.1), complex problem reasoning evaluation (Sec4.2), and ablation studies (Sec4.3, 4.4). First, the applicability evaluation tested SoT’s performance across LLMs of different scales, covering mathematical, planning, factual, and logical reasoning, thereby verifying its generalization and stability. Second, the complex problem reasoning evaluation focused on comparing SoT’s superiority in tackling complex tasks on the GSM8K and MATH datasets. Finally, the ablation studies analyzed key components, such as Betti numbers and temperature parameters, to examine their impact on the complexity, expressiveness, and robustness of the reasoning chain. These experiments comprehensively validate the theoretical and practical value of SoT.
4.1. Applicability Evaluation
The reasoning performance of chain-of-thought methods (such as CoT and CoT-SC) varies significantly across models of different sizes, especially on lightweight models. We aim to validate whether SoT can consistently enhance the reasoning capabilities of various LLMs and demonstrate its generalization and stability across models and task scenarios. For this purpose, we applied CoT, CoT-SC (n=5), and SoT to each model, with the default Betti number set to 7 and mapping set to 3, recording the accuracy across nine datasets.
Experimental Configuration: The test model was GPT-4o-mini, with the Betti number set to 7 and mapping set to 3. We selected five different scales of LLMs, including GPT-4o-mini (lightweight), Qwen2.5-Coder-7B-Instruct (7B), Gemma-3-12b-it (12B), Gemma-3-27b-it (27B), and Qwen2.5-VL-72B-Instruct (72B), covering models from lightweight to large-scale. The baseline models were standard CoT and CoT-SC (n=5). The test datasets included GSM8K, SVAMP, MultiArith, ASDiv, AQUA (mathematical reasoning), MMLU (generalization), BBH (multi-task answering), Data (date prediction), and CLUTRR (logical reasoning), covering nine datasets in total.
Experimental Results: The results are shown in Table 1. SoT significantly outperforms the baselines across all models and datasets. For example, on GSM8K with GPT-4o-mini, SoT achieved 96.0%, improving by 10.9% over CoT (85.1%) and 5.9% over CoT-SC (90.1%). On the logical reasoning task CLUTRR, SoT achieved 75.7%, surpassing CoT (65.9%) and CoT-SC (72.4%). On Qwen2.5-VL-72B-Instruct, SoT achieved 89.4% on AQUA, outperforming CoT (81.1%) and CoT-SC (83.9%). For the lightweight model Qwen2.5-Coder-7B-Instruct, SoT improved GSM8K accuracy to 89.1%, 11.9% higher than CoT (77.2%), showing its superiority in resource-constrained scenarios. Furthermore, on Gemma-3-12b-it, SoT achieved 92.5% on SVAMP, close to the 95.8% achieved by Gemma-3-27b-it, indicating that its generalization ability is not limited by model scale. This result demonstrates the significant advantage of SoT’s structured design in improving reasoning performance and stability across models.
4.2. Complex Reasoning Evaluation
To compare the complex problem reasoning capabilities of SoT with mainstream methods, we evaluated SoT and compared it with extensive open-source data from various papers (Chen et al., 2025) . Specifically, we applied the SoT reasoning chain to GPT-4o-mini and tested its reasoning accuracy on the GSM8K and MATH datasets, while comparing it with other methods from the literature.
Experimental Configuration: The test model was GPT-4o-mini, with the Betti number set to 7 and mapping set to 3. The baselines included CoT and various popular chain-of-thought variants (such as MathPrompter, QuaSAR, MathDivide, etc.). The datasets used were GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021).
Method | Model | GSM8k | MATH |
No-CoT (Deng et al., 2024) | Mistral-7B | 38.0% | - |
ICoT-SI (Deng et al., 2024) | Mistral-7B | 51.0% | - |
- | RecurrentBlock-3.5B | 42.1% | - |
MathCoder-CL (Wang et al., 2023b) | Code-Llama-7B | 67.8% | 30.2% |
MAmmoTH (Yue et al., 2023) | Code-Llama-7B | 59.4% | - |
Brain (Chen et al., 2024a) | Code-Llama-7B | 74.0% | - |
SQ-VAE (Wang et al., 2023a) | Llama-2-7B | 40.0% | 7.0% |
Self-Rewarding (Chen et al., 2024b) | Llama-2-7B | 40.0% | 10.7% |
STaR (Zelikman et al., 2022) | Llama-2-7B | 58.2% | 16.0% |
ENVISIONS (Xu et al., 2024) | Llama-2-7B | 59.0% | 19.0% |
MetaMath (Yu et al., 2023b) | Llama-2-7B | 66.5% | - |
ToRA-Code (Gou et al., 2023) | Llama-2-7B | 72.6% | - |
OVM (Yu et al., 2023a) | Llama-2-7B | 73.7% | - |
- | Llama-3.1-8B | 56.7% | 20.3% |
- | Llama-3.1-70B | 85.5% | 41.4% |
- | Llama-3.1-405B | 89.0% | 53.8% |
- | NuminaMath-7B-CoT | 75.4% | 55.2% |
- | DeepSeek-Coder-7B | 77.4% | 44.4% |
- | Qwen2-7B | 79.9% | 44.2% |
- | Qwen2-Math-7B | 80.4% | 50.4% |
SIaM (Yu et al., 2024) | Qwen-2-Math-Base | 81.5% | 50.0% |
- | Internlm2-math-plus-7B | 84.0% | 54.4% |
OMI2 (Li et al., 2025) | Qwen2.5-Coder-7B | 84.1% | 72.3% |
CODEI/O++ (Li et al., 2025) | Qwen2.5-Coder-7B | 85.7% | 72.1% |
PyEdu (Li et al., 2025) | Qwen2.5-Coder-7B | 85.8% | 71.4% |
CODEI/O (Li et al., 2025) | Qwen2.5-Coder-7B | 86.4% | 71.9% |
OC-SFT-1 (Li et al., 2025) | Qwen2.5-Coder-7B | 86.7% | 70.9% |
WI (Li et al., 2025) | Qwen2.5-Coder-7B | 87.0% | 71.4% |
WI (Full) (Li et al., 2025) | Qwen2.5-Coder-7B | 87.0% | 71.1% |
OMI2 (Full) (Li et al., 2025) | Qwen2.5-Coder-7B | 88.5% | 73.2% |
- | DeepSeekMath-7B-RL | 88.2% | 51.7% |
CoMAT (Leang et al., 2024) | GPT-4 | 93.7% | - |
CoT (Ranaldi et al., 2025) | GPT-4 | 94.5% | - |
FCoT (Lyu et al., 2023) | GPT-4 | 95.0% | - |
MathPrompter (Imani et al., 2023) | GPT-4 | 95.6% | - |
QuaSAR (Radford et al., 2018) | GPT-4 | 96.5% | - |
MathDivide (Srivastava and Gandhi, 2024) | GPT-4 | 96.8% | - |
SoT (Ours) | GPT-4o-mini | 96.0% | 79.1% |
Experimental Results: As shown in Table 2, SoT achieved 96.0% on GSM8K and 79.1% on MATH on GPT-4o-mini. On GSM8K, SoT’s performance is close to the best result achieved by GPT-4 (e.g., MathDivide’s 96.8%) (Chen et al., 2025) and outperforms the best 7B model (e.g., OMI2 Full’s 88.5%) (Chen et al., 2025). On MATH, SoT’s 79.1% significantly outperforms the best mainstream 7B models (e.g., OMI2 Full’s 73.2%). This result indicates that SoT can achieve reasoning capabilities close to those of closed-source large models on lightweight models, significantly narrowing the performance gap between open-source models and GPT-4, validating its superiority in complex mathematical reasoning.
4.3. Betti Number Sensitivity
The SoT reasoning chain regulates topological constraints through the Betti number, and its value directly affects performance. To investigate the optimal configuration of the Betti number and its dynamic impact on structural expressiveness and regularization effects, we designed ablation experiments to reveal its key role in optimizing the reasoning process. For this, we systematically adjusted the Betti number and recorded accuracy changes under different mapping configurations. The goal of the experiment was to analyze the impact of the Betti number on the reasoning chain’s expressiveness and identify the performance saturation point. The results are presented in Figure 5, showing the trend of accuracy changes with varying Betti numbers, with particular attention to its non-monotonic properties and optimal configuration.
Experimental Configuration: The experiment was based on GPT-4o-mini and conducted on the GSM8K dataset. The Betti number was incremented from 1 to 10, with mapping set to 1, 2, and 3 to evaluate performance under different topological decomposition granularities.
Experimental Results: As shown in Figure 5, accuracy changed non-monotonically as the Betti number increased under mapping configurations of 1, 2, and 3. Specifically, when the Betti number started from 1, accuracy improved significantly, indicating that a small number of topological constraints effectively enhanced structural expressiveness. However, as the Betti number increased further, the performance gain gradually diminished and stabilized. The optimal saturation point was reached at 7, beyond which there was no noticeable improvement. The results reveal the dynamic regulatory effect of the Betti number in reasoning chain modeling, and that moderate topological constraints are crucial for optimizing SoT’s structured design, providing strong support for its theoretical foundation.
4.4. Stability Analysis
The temperature parameter influences the diversity of content generated by LLMs, which may challenge reasoning stability. We aimed to explore whether the SoT method could maintain its performance advantage under varying temperature conditions, thereby validating the adaptability of its structured design. To this end, we systematically evaluated SoT and CoT under different temperature conditions, focusing on the stability of accuracy and the relationship between reasoning diversity and consistency. The experiment involved adjusting the temperature step by step and observing the performance trends of the two methods under changing diversity.
Experimental Configuration: The experiment was based on GPT-4o-mini, with the Betti number fixed at 7 and mapping set to 3. The nine test datasets included GSM8K, SVAMP, MultiArith, ASDiv, AQUA (mathematical reasoning), MMLU (generalization), BBH, Date (multi-task QA), and CLUTRR (logical reasoning). The temperature parameter was adjusted from 0.0 to 1.0 in steps of 0.1, and the performance of SoT and CoT was compared. Five reasoning paths were generated in each experiment, and the average accuracy was recorded to reduce the impact of randomness.
Experimental Results:As shown in Figure 6, the accuracy of SoT across multiple datasets varies minimally with temperature (from 0.0 to 1.0), exhibiting stability with small amplitude. Figure 8indicates that the accuracy of CoT fluctuates significantly within the same temperature range, with the amplitude increasing as the temperature rises, especially at high temperatures (close to 1.0), where the fluctuations become more pronounced. The boxplot in Figure 7further compares the accuracy distributions of both methods across datasets: the box of SoT is relatively small, with few outliers, and its accuracy is concentrated with high consistency across datasets, even at high temperatures, without significant outliers. In contrast, the box of CoT is larger, with a more scattered distribution and more outliers, especially at high temperatures, where the reasoning consistency diminishes. Analysis suggests that the high temperature-induced diversity in CoT weakens logical coherence, while SoT maintains reasoning stability.
5. conclusion
In this paper, we introduce the SoT reasoning framework, integrating the MFR principle from algebraic geometry and homological algebra into the CoT paradigm, significantly enhancing LLM performance in complex reasoning tasks. SoT addresses traditional CoT limitations in high-dimensional, nonlinear, and logical scenarios by employing a structured decomposition strategy—utilizing Module, Freeness, Mapping, Exactness, Minimality, and Betti numbers—to create compact, logically consistent reasoning units.
While SoT has shown excellent reasoning capabilities with LLMs, further exploration is needed. We plan to extend the topological decomposition to multimodal reasoning, including inputs like images and tables, to verify its adaptability in cross-modal tasks. Additionally, incorporating iterative MFR concepts will enhance step-by-step problem refinement, optimizing reasoning paths dynamically and improving SoT’s efficiency and accuracy in multimodal and high-dimensional tasks.

### v3

@@
-![[Uncaptioned image]](extracted/6365550/Fig/teaser.png) Syzygy of Thoughts: Improving LLM CoT with the Minimal Free Resolution
-Abstract.
-Chain-of-Thought (CoT) prompting enhances the reasoning of large language models (LLMs) by decomposing problems into sequential steps, mimicking human logic and reducing errors.
-However, complex tasks with vast solution spaces and vague constraints often exceed the capacity of a single reasoning chain.
-Inspired by Minimal Free Resolution (MFR) in commutative algebra and algebraic geometry, we propose Syzygy of Thoughts (SoT)—a novel framework that extends CoT by introducing auxiliary, interrelated reasoning paths.
-SoT captures deeper logical dependencies, enabling more robust and structured problem-solving.
-MFR decomposes a module into a sequence of free modules with minimal rank, providing a structured analytical approach to complex systems.
-This method introduces the concepts of ”Module”, ”Betti numbers”,”Freeness”, ”Mapping”, ”Exactness” and ”Minimality”, enabling the systematic decomposition of the original complex problem into logically complete minimal subproblems while preserving key problem features and reducing reasoning length.
-We tested SoT across diverse datasets (e.g., GSM8K, MATH) and models (e.g., GPT-4o-mini, Qwen2.5), achieving inference accuracy that matches or surpasses mainstream CoTs standards.
-Additionally, by aligning the sampling process with algebraic constraints, our approach enhances the scalability of inference time in LLMs, ensuring both transparent reasoning and high performance.
-Our code will be publicly available at https://github.com/dlMARiA/Syzygy-of-thoughts
-1. Introduction
-The rapid advancements in natural language processing (Vaswani et al., 2017; Devlin et al., 2019; Radford et al., 2018; Brown et al., 2020) have demonstrated outstanding reasoning capabilities across various tasks (Zhao et al., 2023b; Min et al., 2023).
-In recent years, the Chain of Thought (CoT) prompting method (Wei et al., 2022) has emerged as an effective strategy for simulating human reasoning by generating intermediate reasoning steps, significantly enhancing model performance in solving complex problems (Chen et al., 2025; Qu et al., 2025; Sui et al., 2025; Wei et al., 2025; Xia et al., 2024; Liu et al., 2025).
+Syzygy of Thoughts: Improving LLM CoT with the Minimal Free Resolution
+Abstract
+Chain-of-Thought prompting enhances the reasoning of large language models by decomposing problems into sequential steps, mimicking human logic and reducing errors.
+However, for complex tasks with vast solution spaces and ambiguous constraints, a single reasoning chain often proves insufficient.
+Inspired by Minimal Free Resolutions (MFR) in commutative algebra and algebraic geometry, we propose Syzygy of Thoughts (SoT)—a new framework that extends CoT by introducing auxiliary, interrelated reasoning paths.
+By explicitly modeling dependencies among these paths, SoT captures deeper logical structure and enables more robust, structured problem solving.
+MFR decomposes a module into a sequence of free modules of minimal rank, providing a principled way to analyze complex systems.
+Drawing on MFR’s principle of “minimal representation, precise constraints, and layer-by-layer relation tracing,” SoT systematically decomposes an original complex problem into logically complete minimal subproblems while preserving key problem characteristics and reducing reasoning length.
+We evaluate SoT across diverse datasets (e.g., GSM8K, MATH) and model families (e.g., GPT-4o-mini, Qwen2.5).
+Results show that SoT achieves inference accuracy comparable to or exceeding mainstream CoT baselines.
+Moreover, by aligning the sampling process with algebraic constraints, SoT improves the scalability of inference-time computation, offering both transparent reasoning and strong performance.
+The code is available at https://github.com/dlMARiA/Syzygy-of-thoughts
+Syzygy of Thoughts: Improving LLM CoT with the Minimal Free Resolution
+Chenghao Li1, Chaoning Zhang1††thanks: Corresponding author, Yi Lu2, Jiaquan Zhang1, Qigan Sun3, Xudong Wang3, Jiwei Wei1, Guoqing Wang1, Yang Yang1, Heng Tao Shen4,1 1UESTC; 2CNU; 3KHU; 4Tongji Univ.
+1 Introduction
+The rapid advancements in natural language processing Vaswani et al.
+(2017); Devlin et al.
+(2019); Radford et al.
+(2018); Brown et al.
+(2020) have demonstrated outstanding reasoning capabilities across various tasks Zhao et al.
+(2023b); Min et al.
+(2023).
+In recent years, the Chain of Thought (CoT) prompting method Wei et al.
+(2022) has emerged as an effective strategy for simulating human reasoning by generating intermediate reasoning steps, significantly enhancing model performance in solving complex problems.
 The core advantage of the CoT framework lies in its ability to decompose problems into a series of explicit and easily interpretable intermediate steps, thereby improving the transparency and explainability of the reasoning process.
 This explicit stepwise reasoning not only facilitates in-depth analysis and debugging of model outputs but also provides a clear logical pathway for solving complex problems.
-However, despite its success in many tasks, CoT exhibits inherent limitations when dealing with high-dimensional, nonlinear, and abstract logical problems (Wang et al., 2022a; Shi et al., 2023; Jin et al., 2024).
-First, traditional CoT methods often rely on fixed or heuristic decomposition strategies, which struggle to capture essential details and intrinsic connections in multidimensional problems (Wang et al., 2022a).
-Second, due to the lack of systematic structural planning in intermediate steps, the reasoning process tends to include redundant computations and repetitive paths, leading to increased computational resource consumption and inference latency (Jin et al., 2024).
-Although existing improvements have alleviated these issues to some extent, introducing techniques such as multi-path exploration, voting mechanisms, and hierarchical decomposition strategies to enhance robustness (Yao et al., 2023a; Besta et al., 2024; Wang et al., 2022b; Zhou et al., 2022; Yao et al., 2023b), however, they have yet to fully address logical incompleteness and the inadequate decomposition of high-dimensional information.
-As a result, when handling highly complex reasoning tasks, model robustness and accuracy still face significant challenges.
-To address these challenges, we propose a novel reasoning framework—Syzygy of Thoughts (SoT).
-This method draws inspiration from the concept of MFR in algebraic geometry and computational algebra (Eisenbud, 2013; Rossi and Sharifan, 2009; Botbol et al., 2021), systematically decomposing and reconstructing problems by constructing minimal free module sequences.
+However, despite its success in many tasks, CoT faces inherent limitations with high-dimensional, nonlinear, and abstract logical problems Wang et al.
+(2022a); Shi et al.
+(2023); Jin et al.
+(2024).
+Traditional CoT methods rely on fixed or heuristic decompositions, which fail to capture essential details in multidimensional problems Wang et al.
+(2022a).
+Additionally, the lack of structured planning in intermediate steps leads to redundant computations and increased inference latency Jin et al.
+(2024).
+While improvements like multi-path exploration, voting, and hierarchical strategies have helped, they still fail to fully address logical incompleteness and poor high-dimensional decomposition.
+Thus, challenges in robustness and accuracy remain for complex reasoning tasks.
+To address these challenges, we propose a new reasoning framework—Syzygy of Thoughts (SoT).
+“Syzygy” is derived from the Greek word syzygia (), meaning “union” or “pairing”.
+This method draws inspiration from the concept of MFR in algebraic geometry and computational algebra Eisenbud (2013); Rossi and Sharifan (2009); Botbol et al.
+(2021), systematically decomposing and reconstructing problems by constructing minimal free module sequences.
 The incorporation of MFR theory ensures a more rigorous and efficient problem decomposition process while preserving the core characteristics of the problem, effectively reducing redundant computations and logical inconsistencies.
-Specifically, SoT retains the stepwise reasoning advantages of CoT while introducing algebraic tools such as ”module”, ”Betti numbers”, ”freeness”, ”mapping”, ”exactness” and ”minimality” to deeply explore and systematically express the intrinsic structure of complex problems.
-This approach provides a novel theoretical perspective for analyzing high-dimensional and multivariable problems.
+Specifically, SoT retains the stepwise reasoning advantages of CoT while introducing algebraic tools such as "module", "Betti numbers", "freeness", "mapping", "exactness" and "minimality" to deeply explore and systematically express the intrinsic structure of complex problems.
+This approach provides a new theoretical perspective for analyzing high-dimensional and multivariable problems.
 To validate the effectiveness of SoT, we conducted extensive experiments across diverse datasets, including GSM8K, MATH, and others, spanning mathematical reasoning and high-dimensional problem domains.
@@
 Experimental results show that SoT achieves inference accuracy that matches or surpasses mainstream CoTs standards, demonstrating its superior performance and practical potential.
-Finally, the SoT framework exhibits multiple advantages, including an intuitive and transparent reasoning path and improved interpretability of intermediate information.
-The main contributions of this work are as follows:
-Pioneering the integration of MFR into LLM reasoning, enabling structured decomposition of complex problems.
-Developing a novel framework that builds on CoT’s strengths while substantially improving reasoning clarity and structural efficiency.
-Providing theoretical insights and empirical evidence that SoT significantly enhances performance in high-dimensional reasoning tasks.
-2. Related Work
-2.1. Advanced Reasoning Techniques in LLMs
-The reasoning capabilities of LLMs have advanced significantly with breakthroughs in prompting paradigms, notably CoT prompting proposed by Wei et al.(Wei et al., 2022).
-CoT decomposes complex problems into stepwise solutions, mimicking human reasoning, yet struggles with multidimensional or nonlinear tasks like abstract logic and advanced mathematics.
-To address these limitations, variants such as Zero-shot-CoT(Kojima et al., 2022), Self-Consistency CoT (Wang et al., 2022b), and Auto-CoT (Zhang et al., 2022) have emerged, alongside verification-based approaches like VerifyCoT (Zhao et al., 2023a) and CoF-CoT (Nguyen et al., 2023), yielding measurable improvements.
-Structured prompting strategies further enhance reasoning.
-Least-to-Most Prompting (LtM)(Zhou et al., 2022) breaks tasks into manageable subproblems, reducing single-step errors.
-Programmatic methods, including Program of Thought (PoT)(Chen et al., 2022), Chain of Code (CoC)(Li et al., 2023), and Buffer of Thought (BoT)(Yang et al., 2025), integrate discrete variables and procedural steps.
-Building on these, Algorithm of Thought (AoT) (Sel et al., 2023) iteratively synthesizes and refines solutions, boosting token efficiency.
-Tree- and graph-based paradigms offer alternative exploration.
-Tree-of-Thought (ToT)(Yao et al., 2023a) employs a hierarchical structure to explore multiple decision paths, though exponential growth limits deep efficiency.
-Graph-of-Thought (GoT)(Besta et al., 2024) extends this with flexible graph configurations, enabling path recalibration and information aggregation.
-Meanwhile, Skeleton-of-Thought (SoT)(Ning et al., 2023) generates high-level outlines before detailing, balancing speed and quality, while recursive methods like Self-Refine(Madaan et al., 2023) and Step-Back Prompting (Zheng et al., 2023) refine outputs iteratively for higher accuracy.
-2.2. Minimal Free Resolution
-MFR is a core tool in homological algebra (Fieldsteel and Nagel, 2021) and algebraic geometry (Botbol et al., 2021), used to analyze the algebraic structure of modules (Eisenbud, 2013; Rossi and Sharifan, 2009), revealing properties such as rank, symmetry, and relationships (Chen, 2011).
-In computational algebraic geometry, MFR analyzes singularities and invariants of algebraic varieties through syzygies (Evans and Griffith, 1981), optimizing the complexity and efficiency of Gröbner basis computations (Eisenbud et al., 2001; Stewart, 1993; Weispfenning, 1992; Capani et al., 1997).
-In Topological Data Analysis (TDA), MFR accelerates persistent homology computations (Wasserman, 2018), enhancing the stability of topological analysis for high-dimensional data.
-In physics and bioinformatics, MFR is used to analyze Calabi-Yau manifold singularities (Polchinski, 1994), gauge field structures (Hitchin, 2003), and the modular properties of gene regulatory networks (Li et al., 2011).
-The advantages of MFR in optimizing computational complexity and revealing algebraic structures provide a new perspective for its integration into the study of CoT in LLM.
-In CoT, structured reasoning steps are required, but often face challenges of inefficiency and redundancy.
-We innovatively introduce MFR into CoT by decomposing the symbolic dependencies between reasoning steps, eliminating redundant calculations, optimizing intermediate processes, and enhancing transparency.
-This approach opens up a more efficient and structured technical pathway for LLM applications in solving complex problems.
-3. Method
-3.1. Preliminaries
-In algebraic geometry and homological algebra, minimal free resolution is an essential tool for studying the structure and properties of modules, especially when dealing with finitely generated modules over a ring.
-Key Properties: In algebraic geometry and homological algebra, a ”syzygy” refers to the relations among the relations that generate a module.
-A syzygy is a relation among the generators.
-For example, if a module is defined by generators , then a syzygy is a relation of the form:
-where .
-These relations themselves can form a module, which is called the ”first syzygy module.” A module over a ring is an algebraic structure similar to a vector space, but the scalars come from instead of a field.
-A module is called free if it has a basis, i.e., for some dimension (finite or infinite).
-Free modules are analogous to free vector spaces in linear algebra.
-A module is finitely generated if there exists a finite set such that .
-A MFR of a finitely generated module over is an exact sequence:
-Where each is free (), the matrices of the maps contain no units, and the resolution is ”minimal,” meaning the number of free generators is the smallest possible.
-This resolution provides a standardized way to describe the structure of , reflecting both its complexity and its homological properties.
-A MFR is characterized by its minimality, where each uses the smallest number of generators, and the matrices in the maps contain no units.
-It is unique up to isomorphism, meaning any two MFRs of are naturally isomorphic.
-The -th Betti number of , defined as , counts the number of free generators in and reflects the complexity of the resolution.
-3.2. Minimal Free Resolution enhanced CoT
-As shown in Figure 3, the SoT reasoning paradigm employs the MFR framework to model the structure of CoT reasoning paths in LLMs.
+The main contributions of this work:
+We introduce SoT, an MFR-inspired inference-time reasoning framework that generalizes CoT from a single linear chain to interdependent, syzygy-structured reasoning paths.
+We formalize the reasoning process through an MFR lens by mapping algebraic notions (Module, Betti numbers, Freeness, Mapping, Exactness, Minimality) to concrete prompting and search operations, enabling structured decomposition, logical-closure verification, and redundancy pruning.
+We propose an algebraically constrained sampling/selection procedure that improves the accuracy–cost trade-off and provides stable inference under sampling randomness.
+We conduct extensive experiments across nine benchmarks and multiple LLM backbones, showing that SoT consistently outperforms CoT, CoT-SC, GoT, and AoT, with lower or comparable variance, supported by cost and ablation analyses.
+2 Related Work
+2.1 Advanced Reasoning Techniques in LLMs
+Recent advances in prompting have substantially improved the reasoning of LLMs, with Chain-of-Thought (CoT) prompting Wei et al.
+(2022) being a seminal example.
+CoT elicits step-by-step solutions but can still struggle on multidimensional or nonlinear tasks (e.g., abstract logic and advanced mathematics).
+To mitigate these issues, extensions such as Zero-shot-CoT Kojima et al.
+(2022), Self-Consistency CoT Wang et al.
+(2022b), and Auto-CoT Zhang et al.
+(2022) have been proposed, together with verification-based methods including VerifyCoT Zhao et al.
+(2023a) and CoF-CoT Nguyen et al.
+(2023).
+Beyond linear CoT, structured prompting further strengthens reasoning.
+Least-to-Most Prompting (LtM) Zhou et al.
+(2022) decomposes tasks into simpler subproblems, while programmatic paradigms—Program of Thought (PoT) Chen et al.
+(2022), Chain of Code (CoC) Li et al.
+(2023), and Buffer of Thought (BoT) Yang et al.
+(2025)—integrate discrete variables and procedural execution.
+Algorithm of Thought (AoT) Sel et al.
+(2023) then iteratively synthesizes and refines intermediate results to improve token efficiency.
+Search-based paradigms explore multiple solution paths.
+Tree-of-Thought (ToT) Yao et al.
+(2023) performs hierarchical branching but may suffer from exponential growth, whereas Graph-of-Thought (GoT) Besta et al.
+(2024) generalizes this to flexible graph structures for aggregation and recalibration.
+Complementary approaches such as Skeleton-of-Thought (SoT) Ning et al.
+(2023), Self-Refine Madaan et al.
+(2023), and Step-Back Prompting Zheng et al.
+(2023) emphasize outlining and iterative refinement to improve accuracy and stability.
+2.2 Minimal Free Resolution
+MFR is a fundamental tool in homological algebra Fieldsteel and Nagel (2021) and algebraic geometry Botbol et al.
+(2021) for characterizing module structures Eisenbud (2013); Rossi and Sharifan (2009) and deriving invariants such as rank, symmetry, and relations Chen (2011).
+In computational algebraic geometry, syzygy-based MFR supports the study of singularities and invariants of varieties Evans and Griffith (1981) and improves the complexity of Gröbner-basis computations Eisenbud et al.
+(2001); Stewart (1993); Weispfenning (1992); Capani et al.
+(1997).
+Beyond pure algebra, MFR has been used to speed up persistent-homology computation in TDA Wasserman (2018) and to analyze structures arising in physics and bioinformatics, including Calabi–Yau singularities Polchinski (1994), gauge fields Hitchin (2003), and gene regulatory networks Li et al.
+(2011).
+These strengths—exposing structured dependencies while reducing computational redundancy—motivate our integration of MFR into CoT for LLM reasoning.
+By decomposing symbolic dependencies among intermediate steps, we prune redundant branches, streamline intermediate computation, and improve the transparency and efficiency of reasoning, providing a more structured pathway for solving complex problems.
+3 Method
+3.1 Preliminaries
+In algebraic geometry and homological algebra, Minimal Free Resolution is a foundational tool for analyzing finitely generated modules over a ring, offering a structured way to reveal their internal dependencies and complexity.
+Key Concepts. A module over a ring generalizes vector spaces by allowing scalars from . is free if it admits a basis: ; it is finitely generated if a finite set spans via .
+A syzygy refers to a relation among generators, such as:
+and the collection of all such relations forms the first syzygy module.
+An MFR of is an exact sequence:
+where each is free and the maps contain no unit elements. The resolution is minimal if the number of generators in each is smallest possible, making the structure unique up to isomorphism.
+The -th Betti number, , quantifies the number of generators at each level and reflects the complexity of .
+Motivation for SoT.
+As illustrated in Figure 3, we reinterpret CoT reasoning within LLMs through the lens of MFR.
+Each reasoning stage—e.g., problem understanding, decomposition, subgoal chaining, and final decision—is treated as a generator or relation within a module.
+This enables SoT to capture latent logical dependencies and hierarchical structure, moving beyond CoT’s linear stepwise form toward a more principled and algebraically grounded representation.
+The SoT reasoning paradigm employs the MFR framework to model the structure of CoT reasoning paths in LLMs.
 Each reasoning step—such as problem interpretation, task decomposition, subgoal chaining, and final inference—is abstracted as generators or relations within modules, reinterpreting CoT as algebraic objects with generative structures and latent dependencies, rather than merely a flat sequence of symbols.
-![[Uncaptioned image]](extracted/6365550/Fig/modul.png)
 Module: Initial Complex Problem
@@
 This stage ensures that downstream processes operate within well-defined boundaries, setting the stage for structured analysis.
-![[Uncaptioned image]](extracted/6365550/Fig/betti.png)
 Betti Numbers: Quantifying Complexity
@@
 Betti numbers thus guide reasoning efficiency, balancing structural complexity with computational simplicity.
-![[Uncaptioned image]](extracted/6365550/Fig/freeness.png)
 Freeness
@@
 This reduces dimensionality and provides clear entry points for resolution, enhancing the tractability of subsequent steps while maintaining logical coherence.
-![[Uncaptioned image]](extracted/6365550/Fig/mapping.png)
 Mapping: Solving Strategies
 Effective problem-solving requires systematic strategies that map auxiliary conditions to actionable reasoning paths.
-These paths must exhibit (i) directness, minimizing redundant operations, and (ii) logical soundness, grounding each step in explicit premises).
+These paths must exhibit (i) directness, minimizing redundant operations, and (ii) logical soundness, grounding each step in explicit premises.
 LLMs iteratively solve subproblems along these paths, linking current conditions to new conclusions, ensuring continuity and interpretability throughout the reasoning chain.
-![[Uncaptioned image]](extracted/6365550/Fig/exactness.png)
 Exactness: Logical Completeness
@@
 This ensures a coherent and reliable framework, minimizing errors from incomplete logic.
-![[Uncaptioned image]](extracted/6365550/Fig/mini.png)
 Minimality: Optimal Solutions
@@
 This reduces overall complexity, enhancing both computational efficiency and solution transparency.
-After introducing the above six concepts, LLMs can be directly stimulated in a conversational manner.
+After introducing the above six concepts, LLMs can be directly stimulated conversationally.
 Figure 6 illustrates the conversational framework of these six concepts.
@@
 Finally, minimality and accuracy are used to select the optimal solution.
-3.3. Syzygy of Thoughts
+3.2 Syzygy of Thoughts
 The process of LLM reasoning for complex problems can be viewed as solving a complex reasoning problem through MFR.
-The entire process is formalized via a MFR sequence, where the structured representation of the reasoning process is a set of mappings induced by the LLM, forming the following sequence:
+The entire process is formalized via an MFR sequence, where the structured representation of the reasoning process is a set of mappings induced by the LLM, forming the following sequence:
 Moreover, we define the intermediate space as a free module constructed from basic reasoning units:
-where:
-- is the reasoning space (LLM Reasoning Space), - are the basic units (Tokens) used in the reasoning construction, - Each is induced by the LLM, representing a single-step logical deduction in the CoT.
-3.3.1. SoT Reasoning.
-The Algorithm 1 and Figure 3 explain the entire process of SoT.
+where is the reasoning space (LLM Reasoning Space), are the basic units (Tokens) used in the reasoning construction.
+Each is induced by the LLM, representing a single-step logical deduction in the CoT.
+3.2.1 SoT Reasoning.
+Figure 3 explains the entire process of SoT.
 The SoT framework begins with an initial complex problem , characterized by high-dimensional structure, intricate logical dependencies, and ambiguous constraint relationships.
@@
 Finally, the optimal reasoning sequence is parsed using structured pattern-matching strategies (e.g., regular expressions, JSON parsing) to extract the final solution.
-Method | Math Reasoning | Gene. | Multitask | Temporal | Log. | ||||
+| Method | Math Reasoning | Gene. | Mul. | Tem. | Log. | ||||
 | GSM8K | SVAMP | MultiArith | ASDiv | AQUA | MMLU | BBH | Date | CLUTRR | |
 | GPT-4o-mini | |||||||||
-CoT | 85.1% | 84.4% | 99.2% | 97.0% | 65.0% | 63.1% | 66.3% | 51.8% | 65.9% |
-CoT-SC (n=5) | 90.1% | 86.0% | 99.5% | 98.5% | 70.9% | 67.3% | 69.2% | 54.9% | 72.4% |
-SoT (Ours) | 96.0% | 92.2% | 99.7% | 99.8% | 75.6% | 75.2% | 72.8% | 75.2% | 75.7% |
+| CoT (Wei et al., 2022) | 85.4% | 84.7% | 89.5% | 92.3% | 65.3% | 63.4% | 66.6% | 52.1% | 66.2% |
+| CoT-SC (Wang et al., 2022b) | 89.9% | 85.8% | 89.7% | 93.3% | 70.7% | 67.1% | 69.0% | 54.7% | 72.2% |
+| GoT (Besta et al., 2024) | 93.2% | 89.2% | 90.9% | 94.2% | 73.3% | 71.3% | 71.1% | 65.2% | 74.2% |
+| AoT (Teng et al., 2025) | 95.3% | 91.5% | 91.0% | 94.1% | 75.1% | 74.7% | 72.4% | 74.7% | 75.2% |
+| SoT (Ours) | 96.4% | 92.6% | 92.1% | 95.2% | 76.0% | 75.6% | 73.2% | 75.6% | 76.1% |
 | Qwen2.5-Coder-7B-Instruct | |||||||||
-CoT | 77.2% | 82.4% | 92.3% | 92.0% | 60.6% | 55.1% | 47.1% | 31.0% | 20.1% |
-CoT-SC (n=5) | 80.2% | 84.1% | 95.0% | 95.0% | 62.2% | 56.3% | 49.3% | 32.9% | 21.0% |
-SoT (Ours) | 89.1% | 90.6% | 97.0% | 99.8% | 63.3% | 57.1% | 57.3% | 36.2% | 26.3% |
+| CoT (Wei et al., 2022) | 77.5% | 82.7% | 84.6% | 87.3% | 60.9% | 55.4% | 47.4% | 31.3% | 20.4% |
+| CoT-SC (Wang et al., 2022b) | 80.0% | 83.9% | 86.8% | 89.8% | 62.0% | 56.1% | 49.1% | 32.7% | 20.8% |
+| GoT (Besta et al., 2024) | 84.8% | 87.5% | 88.1% | 92.5% | 62.9% | 56.8% | 53.4% | 34.7% | 23.8% |
+| AoT (Teng et al., 2025) | 88.5% | 90.0% | 88.4% | 94.1% | 63.0% | 56.8% | 57.0% | 36.2% | 26.4% |
+| SoT (Ours) | 89.5% | 91.0% | 89.4% | 95.2% | 63.7% | 57.5% | 57.7% | 36.6% | 26.7% |
 | Qwen2.5-VL-72B-Instruct | |||||||||
-CoT | 86.1% | 86.9% | 98.8% | 98.0% | 81.1% | 80.1% | 77.3% | 75.2% | 70.1% |
-CoT-SC (n=5) | 89.1% | 88.2% | 99.3% | 98.4% | 83.9% | 82.9% | 79.0% | 78.0% | 75.0% |
-SoT (Ours) | 96.0% | 95.8% | 99.7% | 99.2% | 89.4% | 84.3% | 85.3% | 80.2% | 78.9% |
+| CoT (Wei et al., 2022) | 86.4% | 87.2% | 90.1% | 92.3% | 81.4% | 80.4% | 77.6% | 75.5% | 70.4% |
+| CoT-SC (Wang et al., 2022b) | 88.9% | 88.0% | 91.1% | 93.6% | 83.7% | 82.7% | 78.8% | 77.8% | 74.8% |
+| GoT (Besta et al., 2024) | 92.7% | 92.1% | 91.6% | 93.1% | 86.8% | 83.7% | 82.2% | 79.2% | 77.0% |
+| AoT (Teng et al., 2025) | 95.3% | 95.1% | 91.0% | 93.5% | 88.8% | 83.7% | 84.7% | 79.7% | 78.4% |
+| SoT (Ours) | 96.4% | 96.2% | 92.1% | 94.6% | 89.8% | 84.7% | 85.7% | 80.6% | 79.3% |
 | Gemma-3-27b-it | |||||||||
-CoT | 83.1% | 85.9% | 91.9% | 98.5% | 80.3% | 70.8% | 70.7% | 76.9% | 65.3% |
-CoT-SC (n=5) | 87.1% | 87.0% | 92.3% | 99.2% | 85.4% | 73.2% | 73.2% | 80.2% | 66.4% |
-SoT (Ours) | 96.0% | 95.8% | 99.7% | 99.2% | 89.4% | 84.3% | 85.3% | 80.2% | 78.9% |
+| CoT (Wei et al., 2022) | 83.4% | 86.2% | 84.2% | 90.8% | 80.6% | 71.1% | 71.0% | 77.2% | 65.6% |
+| CoT-SC (Wang et al., 2022b) | 86.9% | 86.8% | 84.5% | 92.0% | 85.2% | 73.0% | 73.0% | 80.0% | 66.2% |
+| GoT (Besta et al., 2024) | 91.7% | 91.5% | 88.3% | 91.3% | 87.5% | 78.8% | 79.3% | 80.3% | 72.8% |
+| AoT (Teng et al., 2025) | 93.2% | 93.1% | 88.4% | 91.8% | 85.6% | 81.7% | 82.9% | 77.4% | 76.1% |
+| SoT (Ours) | 94.2% | 95.3% | 91.9% | 93.3% | 88.1% | 83.4% | 84.7% | 79.6% | 78.4% |
 | Gemma-3-12b-it | |||||||||
-CoT | 83.2% | 79.0% | 90.4% | 97.7% | 68.9% | 68.1% | 64.6% | 77.7% | 49.0% |
-CoT-SC (n=5) | 86.1% | 81.0% | 93.3% | 98.0% | 71.7% | 70.6% | 66.7% | 80.2% | 52.2% |
-SoT (Ours) | 92.1% | 92.5% | 96.1% | 99.2% | 77.2% | 72.3% | 69.1% | 82.5% | 55.0% |
-4. Experiments
-To evaluate the performance of SoT, we designed a series of experiments, including applicability evaluation (Sec4.1), complex problem reasoning evaluation (Sec4.2), and ablation studies (Sec4.3, 4.4).
-First, the applicability evaluation tested SoT’s performance across LLMs of different scales, covering mathematical, planning, factual, and logical reasoning, thereby verifying its generalization and stability.
-Second, the complex problem reasoning evaluation focused on comparing SoT’s superiority in tackling complex tasks on the GSM8K and MATH datasets.
-Finally, the ablation studies analyzed key components, such as Betti numbers and temperature parameters, to examine their impact on the complexity, expressiveness, and robustness of the reasoning chain.
-These experiments comprehensively validate the theoretical and practical value of SoT.
-4.1. Applicability Evaluation
-The reasoning performance of chain-of-thought methods (such as CoT and CoT-SC) varies significantly across models of different sizes, especially on lightweight models.
-We aim to validate whether SoT can consistently enhance the reasoning capabilities of various LLMs and demonstrate its generalization and stability across models and task scenarios.
-For this purpose, we applied CoT, CoT-SC (n=5), and SoT to each model, with the default Betti number set to 7 and mapping set to 3, recording the accuracy across nine datasets.
-Experimental Configuration: The test model was GPT-4o-mini, with the Betti number set to 7 and mapping set to 3.
-We selected five different scales of LLMs, including GPT-4o-mini (lightweight), Qwen2.5-Coder-7B-Instruct (7B), Gemma-3-12b-it (12B), Gemma-3-27b-it (27B), and Qwen2.5-VL-72B-Instruct (72B), covering models from lightweight to large-scale.
-The baseline models were standard CoT and CoT-SC (n=5).
-The test datasets included GSM8K, SVAMP, MultiArith, ASDiv, AQUA (mathematical reasoning), MMLU (generalization), BBH (multi-task answering), Data (date prediction), and CLUTRR (logical reasoning), covering nine datasets in total.
-Experimental Results: The results are shown in Table 1.
-SoT significantly outperforms the baselines across all models and datasets.
-For example, on GSM8K with GPT-4o-mini, SoT achieved 96.0%, improving by 10.9% over CoT (85.1%) and 5.9% over CoT-SC (90.1%).
-On the logical reasoning task CLUTRR, SoT achieved 75.7%, surpassing CoT (65.9%) and CoT-SC (72.4%).
-On Qwen2.5-VL-72B-Instruct, SoT achieved 89.4% on AQUA, outperforming CoT (81.1%) and CoT-SC (83.9%).
-For the lightweight model Qwen2.5-Coder-7B-Instruct, SoT improved GSM8K accuracy to 89.1%, 11.9% higher than CoT (77.2%), showing its superiority in resource-constrained scenarios.
-Furthermore, on Gemma-3-12b-it, SoT achieved 92.5% on SVAMP, close to the 95.8% achieved by Gemma-3-27b-it, indicating that its generalization ability is not limited by model scale.
-This result demonstrates the significant advantage of SoT’s structured design in improving reasoning performance and stability across models.
-4.2. Complex Reasoning Evaluation
-To compare the complex problem reasoning capabilities of SoT with mainstream methods, we evaluated SoT and compared it with extensive open-source data from various papers (Chen et al., 2025) .
-Specifically, we applied the SoT reasoning chain to GPT-4o-mini and tested its reasoning accuracy on the GSM8K and MATH datasets, while comparing it with other methods from the literature.
-Experimental Configuration: The test model was GPT-4o-mini, with the Betti number set to 7 and mapping set to 3.
-The baselines included CoT and various popular chain-of-thought variants (such as MathPrompter, QuaSAR, MathDivide, etc.).
-The datasets used were GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021).
-Method | Model | GSM8k | MATH |
-No-CoT (Deng et al., 2024) | Mistral-7B | 38.0% | - |
-ICoT-SI (Deng et al., 2024) | Mistral-7B | 51.0% | - |
-- | RecurrentBlock-3.5B | 42.1% | - |
-MathCoder-CL (Wang et al., 2023b) | Code-Llama-7B | 67.8% | 30.2% |
-MAmmoTH (Yue et al., 2023) | Code-Llama-7B | 59.4% | - |
-Brain (Chen et al., 2024a) | Code-Llama-7B | 74.0% | - |
-SQ-VAE (Wang et al., 2023a) | Llama-2-7B | 40.0% | 7.0% |
-Self-Rewarding (Chen et al., 2024b) | Llama-2-7B | 40.0% | 10.7% |
-STaR (Zelikman et al., 2022) | Llama-2-7B | 58.2% | 16.0% |
-ENVISIONS (Xu et al., 2024) | Llama-2-7B | 59.0% | 19.0% |
-MetaMath (Yu et al., 2023b) | Llama-2-7B | 66.5% | - |
-ToRA-Code (Gou et al., 2023) | Llama-2-7B | 72.6% | - |
-OVM (Yu et al., 2023a) | Llama-2-7B | 73.7% | - |
-- | Llama-3.1-8B | 56.7% | 20.3% |
-- | Llama-3.1-70B | 85.5% | 41.4% |
-- | Llama-3.1-405B | 89.0% | 53.8% |
-- | NuminaMath-7B-CoT | 75.4% | 55.2% |
-- | DeepSeek-Coder-7B | 77.4% | 44.4% |
-- | Qwen2-7B | 79.9% | 44.2% |
-- | Qwen2-Math-7B | 80.4% | 50.4% |
-SIaM (Yu et al., 2024) | Qwen-2-Math-Base | 81.5% | 50.0% |
-- | Internlm2-math-plus-7B | 84.0% | 54.4% |
-OMI2 (Li et al., 2025) | Qwen2.5-Coder-7B | 84.1% | 72.3% |
-CODEI/O++ (Li et al., 2025) | Qwen2.5-Coder-7B | 85.7% | 72.1% |
-PyEdu (Li et al., 2025) | Qwen2.5-Coder-7B | 85.8% | 71.4% |
-CODEI/O (Li et al., 2025) | Qwen2.5-Coder-7B | 86.4% | 71.9% |
-OC-SFT-1 (Li et al., 2025) | Qwen2.5-Coder-7B | 86.7% | 70.9% |
-WI (Li et al., 2025) | Qwen2.5-Coder-7B | 87.0% | 71.4% |
-WI (Full) (Li et al., 2025) | Qwen2.5-Coder-7B | 87.0% | 71.1% |
-OMI2 (Full) (Li et al., 2025) | Qwen2.5-Coder-7B | 88.5% | 73.2% |
-- | DeepSeekMath-7B-RL | 88.2% | 51.7% |
-CoMAT (Leang et al., 2024) | GPT-4 | 93.7% | - |
-CoT (Ranaldi et al., 2025) | GPT-4 | 94.5% | - |
-FCoT (Lyu et al., 2023) | GPT-4 | 95.0% | - |
-MathPrompter (Imani et al., 2023) | GPT-4 | 95.6% | - |
-QuaSAR (Radford et al., 2018) | GPT-4 | 96.5% | - |
-MathDivide (Srivastava and Gandhi, 2024) | GPT-4 | 96.8% | - |
-SoT (Ours) | GPT-4o-mini | 96.0% | 79.1% |
-Experimental Results: As shown in Table 2, SoT achieved 96.0% on GSM8K and 79.1% on MATH on GPT-4o-mini.
-On GSM8K, SoT’s performance is close to the best result achieved by GPT-4 (e.g., MathDivide’s 96.8%) (Chen et al., 2025) and outperforms the best 7B model (e.g., OMI2 Full’s 88.5%) (Chen et al., 2025).
-On MATH, SoT’s 79.1% significantly outperforms the best mainstream 7B models (e.g., OMI2 Full’s 73.2%).
-This result indicates that SoT can achieve reasoning capabilities close to those of closed-source large models on lightweight models, significantly narrowing the performance gap between open-source models and GPT-4, validating its superiority in complex mathematical reasoning.
-4.3. Betti Number Sensitivity
+| CoT (Wei et al., 2022) | 83.5% | 79.3% | 82.7% | 91.0% | 69.2% | 68.4% | 64.9% | 78.0% | 49.3% |
+| CoT-SC (Wang et al., 2022b) | 85.9% | 80.8% | 85.1% | 93.2% | 71.5% | 70.4% | 66.5% | 80.0% | 52.0% |
+| GoT (Besta et al., 2024) | 89.2% | 86.8% | 86.8% | 93.9% | 74.5% | 71.6% | 68.0% | 81.5% | 53.7% |
+| AoT (Teng et al., 2025) | 91.4% | 91.8% | 87.5% | 93.5% | 76.7% | 71.9% | 68.7% | 82.0% | 54.8% |
+| SoT (Ours) | 92.5% | 92.9% | 88.5% | 94.6% | 77.6% | 72.7% | 69.5% | 82.9% | 55.4% |
+4 Experiments
+4.1 Applicability Evaluation
+Configuration.
+As shown in Table 1, we compare the proposed SoT method with four representative reasoning strategies: Chain-of-Thought (CoT) (Wei et al., 2022), Self-Consistency CoT (CoT-SC) (Wang et al., 2022b), Graph-of-Thought (GoT) (Besta et al., 2024), and Atom-of-Thought (AoT) (Teng et al., 2025).
+The evaluation covers nine benchmarks spanning four categories: mathematical word problems (GSM8K, SVAMP, MultiArith, ASDiv, AQUA), general-knowledge and multitask QA (MMLU, BBH), temporal reasoning (Date), and relational logical reasoning (CLUTRR).
+For CoT-SC, we follow the standard setting with sampled reasoning paths.
+GoT is instantiated with beam size and search depth , while AoT uses a maximum search depth of , consistent with the original works.
+We instantiate all methods on five backbone LLMs: GPT-4o-mini, Qwen2.5-Coder-7B-Instruct, Qwen2.5-VL-72B-Instruct, Gemma-3-27b-it, and Gemma-3-12b-it.
+For each backbone, we keep decoding hyper-parameters (e.g., temperature, top-, maximum generation length) fixed across methods and only change the reasoning protocol.
+We report mean accuracy and standard deviation over multiple random seeds to account for sampling variance.
+Results.
+From Table 1, SoT consistently achieves the best performance across all benchmarks and all backbone models.
+On the mathematical reasoning benchmarks, SoT improves over the strongest baseline AoT by around percentage point on GSM8K and SVAMP for every model family (e.g., on GPT-4o-mini, GSM8K accuracy increases from to ), while also yielding smaller but steady gains on MultiArith and ASDiv.
+On the more challenging AQUA dataset, SoT brings additional improvements of roughly – points across different backbones, indicating that searching over structured thoughts helps the model better handle multi-step algebraic reasoning and eliminate distractor options.
+For general-knowledge and multitask reasoning, SoT further improves MMLU and BBH by approximately – points over AoT on all models, with Qwen2.5-VL-72B and Gemma-3-27b-it achieving the highest absolute accuracies.
+Finally, on temporal and relational logical reasoning (Date and CLUTRR), SoT again attains the best results under every backbone, with gains of up to about point, suggesting that the structured search procedure is beneficial beyond pure arithmetic reasoning.
+Overall, SoT dominates all baselines in every setting in Table 1 while maintaining comparable or even lower variance, demonstrating that it provides a stable and broadly applicable improvement to LLM reasoning.
+4.2 Cost Test
+To systematically compare the trade-offs between accuracy and search cost across different reasoning frameworks, we construct the three-dimensional bubble plot shown in Figure 4, which places five methods under a unified lens: CoT, CoT-SC, GoT, AoT, and our proposed SoT.
+All methods are evaluated using the same model and the same set of nine benchmark tasks.
+In the plot, the vertical axis represents the macro-average accuracy over the nine tasks, the horizontal axis denotes the average number of candidate reasoning paths generated per problem, and the bubble area is proportional to the average number of tokens consumed per problem.
+We provide the detailed definition of our path-counting metric in Appendix 4.2.
+Based on the above definitions, we visualize a representative set of method configurations.
+Along the horizontal axis, the number of candidate paths increases from left to right; along the vertical axis, the macro-average accuracy improves from bottom to top; and the bubble size grows with the token cost.
+When the number of candidate paths and the token budget are comparable, SoT attains substantially higher macro-average accuracy than CoT-SC, GoT, and AoT.
+Moreover, as the number of syzygies increases from to , the marginal gains diminish noticeably—indicating that SoT achieves near-saturated performance even with only a small number of syzygies.
+Overall, SoT occupies a particularly favorable region in the three-way trade-off space: with a moderate number of candidate paths and an acceptable token cost, it achieves the highest macro-average accuracy and significantly enhances the effectiveness of each individual candidate path through its syzygy-based structure.
+4.3 Abalation Study
+Betti Number Sensitivity.
 The SoT reasoning chain regulates topological constraints through the Betti number, and its value directly affects performance.
@@
 The results are presented in Figure 5, showing the trend of accuracy changes with varying Betti numbers, with particular attention to its non-monotonic properties and optimal configuration.
-Experimental Configuration: The experiment was based on GPT-4o-mini and conducted on the GSM8K dataset.
+Configuration.
+The experiment was based on GPT-4o-mini and conducted on the GSM8K dataset.
 The Betti number was incremented from 1 to 10, with mapping set to 1, 2, and 3 to evaluate performance under different topological decomposition granularities.
-Experimental Results: As shown in Figure 5, accuracy changed non-monotonically as the Betti number increased under mapping configurations of 1, 2, and 3.
+Results.
+As shown in Figure 5 (left), accuracy changed non-monotonically as the Betti number increased under mapping configurations of 1, 2, and 3.
 Specifically, when the Betti number started from 1, accuracy improved significantly, indicating that a small number of topological constraints effectively enhanced structural expressiveness.
@@
 The results reveal the dynamic regulatory effect of the Betti number in reasoning chain modeling, and that moderate topological constraints are crucial for optimizing SoT’s structured design, providing strong support for its theoretical foundation.
-4.4. Stability Analysis
+Stability Analysis.
 The temperature parameter influences the diversity of content generated by LLMs, which may challenge reasoning stability.
@@
 The experiment involved adjusting the temperature step by step and observing the performance trends of the two methods under changing diversity.
-Experimental Configuration: The experiment was based on GPT-4o-mini, with the Betti number fixed at 7 and mapping set to 3.
+Configuration.
+The experiment was based on GPT-4o-mini, with the Betti number fixed at 7 and mapping set to 3.
 The nine test datasets included GSM8K, SVAMP, MultiArith, ASDiv, AQUA (mathematical reasoning), MMLU (generalization), BBH, Date (multi-task QA), and CLUTRR (logical reasoning).
@@
 Five reasoning paths were generated in each experiment, and the average accuracy was recorded to reduce the impact of randomness.
-Experimental Results:As shown in Figure 6, the accuracy of SoT across multiple datasets varies minimally with temperature (from 0.0 to 1.0), exhibiting stability with small amplitude.
-Figure 8indicates that the accuracy of CoT fluctuates significantly within the same temperature range, with the amplitude increasing as the temperature rises, especially at high temperatures (close to 1.0), where the fluctuations become more pronounced.
-The boxplot in Figure 7further compares the accuracy distributions of both methods across datasets: the box of SoT is relatively small, with few outliers, and its accuracy is concentrated with high consistency across datasets, even at high temperatures, without significant outliers.
-In contrast, the box of CoT is larger, with a more scattered distribution and more outliers, especially at high temperatures, where the reasoning consistency diminishes.
-Analysis suggests that the high temperature-induced diversity in CoT weakens logical coherence, while SoT maintains reasoning stability.
-5. conclusion
-In this paper, we introduce the SoT reasoning framework, integrating the MFR principle from algebraic geometry and homological algebra into the CoT paradigm, significantly enhancing LLM performance in complex reasoning tasks.
-SoT addresses traditional CoT limitations in high-dimensional, nonlinear, and logical scenarios by employing a structured decomposition strategy—utilizing Module, Freeness, Mapping, Exactness, Minimality, and Betti numbers—to create compact, logically consistent reasoning units.
-While SoT has shown excellent reasoning capabilities with LLMs, further exploration is needed.
-We plan to extend the topological decomposition to multimodal reasoning, including inputs like images and tables, to verify its adaptability in cross-modal tasks.
-Additionally, incorporating iterative MFR concepts will enhance step-by-step problem refinement, optimizing reasoning paths dynamically and improving SoT’s efficiency and accuracy in multimodal and high-dimensional tasks.
+Results.
+As shown in Figure 5 (right), SoT’s accuracy remains stable across temperature variations (0.0 to 1.0) with minimal fluctuation.
+In contrast, CoT’s accuracy fluctuates significantly, especially at higher temperatures, where the amplitude increases.
+The boxplot in Figure 5 (middle) reveals that SoT’s accuracy is concentrated with little variation and few outliers across datasets, even at high temperatures.
+On the other hand, CoT shows a wider spread with more outliers, especially at higher temperatures, where reasoning consistency decreases.
+This analysis indicates that CoT’s high temperature-induced diversity weakens logical coherence, while SoT maintains stability.
+5 Conclusions
+We introduced SoT, an MFR-inspired test-time reasoning framework that reframes CoT from a single linear trace into a syzygy-structured family of interdependent reasoning paths.
+By mapping MFR notions to concrete prompting and selection operations, SoT explicitly organizes auxiliary conditions, constructs candidate “syzygies,” verifies logical closure, and prunes redundancy to produce compact yet faithful solutions.
+Across nine benchmarks spanning math, general knowledge, multitask QA, temporal reasoning, and relational logic, and across five LLM backbones, SoT consistently outperforms strong baselines while maintaining comparable or lower variance, indicating both improved accuracy and more stable inference.
+Moreover, SoT occupies a favorable region in the accuracy–cost–paths trade-off, achieving near-saturated gains with only a small number of syzygies, and exhibiting strong robustness under sampling diversity.
+6 Limitations
+While SoT demonstrates strong reasoning capabilities, further research is needed.
+We plan to extend the topological decomposition to multimodal reasoning, including images and tables, to test its adaptability in cross-modal tasks.
+Additionally, incorporating iterative MFR concepts will refine problem-solving steps, optimize reasoning paths, and improve SoT’s efficiency in multimodal and high-dimensional tasks.
+Appendix A Appendix
+A.1 Full Pseudocode of the SoT Pipeline
+In this appendix, we present the full pseudocode of the SoT pipeline.
+Algorithm 1 provides an algorithmic implementation of the LLM-based minimal free resolution discussed in Section 3, turning the conceptual modules in Figure 3 and the abstract diagram in Figure 6 into an executable procedure.
+Explanation of Algorithm 1.
+We briefly explain the main components of Algorithm 1, which instantiates the SoT framework described in Section 3.
+Analyse and homological complexity.
+Given a reasoning problem and its metadata , the subroutine Analyse estimates the regularity and the projective dimension .
+These quantities summarize the Betti-number profile introduced in Section 3: determines the number of auxiliary (freeness) conditions, and specifies how many candidate resolutions (syzygies) are explored.
+Larger values of indicate higher problem complexity.
+Generating auxiliary conditions (Freeness).
+The loop over constructs a set of auxiliary conditions, corresponding to the “Freeness” module in Figure 3.
+Each call to LLMGenerateAuxCondition conditions on previously created free modules and maps to propose a new constraint .
+These constraints guide the LLM toward structurally coherent reasoning decompositions.
+Constructing candidate syzygies (Syzygy of Thoughts).
+The loop over implements the “Syzygy of Thoughts” mechanism from Section 3.
+Each call to LLMResolve yields a candidate reasoning chain that satisfies the auxiliary conditions.
+From an algebraic viewpoint, each functions as a syzygy relating different generators of , and the list collects all such candidates.
+Selecting a minimal resolution (Minimality).
+Once the candidate set is formed, the subroutine ArgMinScore selects the best resolution according to a scoring function .
+This function, implemented via an LLM scoring prompt, evaluates both structural regularity (e.g., consistency with freeness and mapping constraints) and task-specific correctness.
+This step aligns with the “Minimality” module in Figure 3.
+Extracting the final answer (Exactness).
+Finally, ExtractFinalAnswer converts the selected reasoning chain into the final task-specific answer.
+Pattern matching techniques (e.g., regular expressions or JSON parsing) ensure the output is consistent with the resolved structure and adheres to benchmark formatting requirements, realizing the “Exactness” principle.
+A.2 Conceptual Illustration of the SoT Pipeline
+Explanation.
+Figure 6 provides a concrete walk-through of the SoT framework using a simple arithmetic problem.
+The illustration makes explicit how SoT transforms a natural-language question into a structured multi-stage reasoning process.
+(1) Module and Betti Numbers.
+The LLM first rewrites the question into a normalized “module” form and performs a lightweight complexity analysis.
+Based on the estimated problem structure, it predicts the required number of auxiliary conditions, corresponding to the Betti-number estimate discussed in Section 3.
+(2) Freeness.
+The LLL generates non-redundant auxiliary conditions (e.g., price per pair, number of children, number of pairs per child).
+These auxiliary facts form the initial free modules , which constrain the search space of valid reasoning chains.
+(3) Mappings.
+The LLM organizes auxiliary conditions into candidate reasoning strategies by combining freeness conditions in different orders.
+Each such combination corresponds to a candidate syzygy, i.e., a structured reasoning chain that reflects different possible mappings between the modules in the minimal free resolution.
+(4) Exactness.
+The LLM inspects each candidate strategy and verifies whether every reasoning step is justified, gap-free, and based on explicitly stated assumptions.
+This matches the exactness requirement in the algebraic SoT formulation.
+(5) Minimality.
+Among all valid strategies, the LLM selects the one with the smallest reasoning footprint, balancing accuracy, complexity, and simplicity.
+This corresponds to the minimality principle described in Section 3.
+(6) Final Answer.
+Once the optimal syzygy is chosen, the LLM extracts the final numerical answer while ensuring consistency with all earlier reasoning stages.
+This mirrors the final extraction step in Algorithm 1.
+Relation to Algorithm 1.
+This example illustrates the same pipeline implemented formally in Algorithm 1: Analyse corresponds to Step 1; LLMGenerateAuxCondition and LLMResolve correspond to Steps 2–3; ArgMinScore corresponds to Step 4; and ExtractFinalAnswer corresponds to Step 5.
+The illustration provides an intuitive view complementary to the algorithmic description.
+A.3 Complex Reasoning Evaluation
+To compare the complex problem reasoning capabilities of SoT with mainstream methods, we evaluated SoT and compared it with extensive open-source data from various papers Chen et al.
+(2025).
+Specifically, we applied the SoT reasoning chain to GPT-4o-mini and tested its reasoning accuracy on the GSM8K and MATH datasets, while comparing it with other methods from the literature.
+Experimental Configuration: The test model was GPT-4o-mini, with the Betti number set to 7 and mapping set to 3.
+The baselines included CoT and various popular chain-of-thought variants (such as MathPrompter, QuaSAR, MathDivide, etc.).
+The datasets used were GSM8K Cobbe et al.
+(2021) and MATH Hendrycks et al.
+(2021).
+| Method | Model | GSM8k | MATH |
+| No-CoT Deng et al. (2024) | Mistral-7B | 38.0% | - |
+| ICoT-SI Deng et al. (2024) | Mistral-7B | 51.0% | - |
+| - | RecurrentBlock-3.5B | 42.1% | - |
+| MathCoder-CL Wang et al. (2023a) | Code-Llama-7B | 67.8% | 30.2% |
+| MAmmoTH Yue et al. (2023) | Code-Llama-7B | 59.4% | - |
+| Brain Chen et al. (2024a) | Code-Llama-7B | 74.0% | - |
+| SQ-VAE Wang et al. (2023b) | Llama-2-7B | 40.0% | 7.0% |
+| Self-Rewarding Chen et al. (2024b) | Llama-2-7B | 40.0% | 10.7% |
+| STaR Zelikman et al. (2022) | Llama-2-7B | 58.2% | 16.0% |
+| ENVISIONS Xu et al. (2024) | Llama-2-7B | 59.0% | 19.0% |
+| MetaMath Yu et al. (2023b) | Llama-2-7B | 66.5% | - |
+| ToRA-Code Gou et al. (2023) | Llama-2-7B | 72.6% | - |
+| OVM Yu et al. (2023a) | Llama-2-7B | 73.7% | - |
+| - | Llama-3.1-8B | 56.7% | 20.3% |
+| - | Llama-3.1-70B | 85.5% | 41.4% |
+| - | Llama-3.1-405B | 89.0% | 53.8% |
+| - | NuminaMath-7B-CoT | 75.4% | 55.2% |
+| - | DeepSeek-Coder-7B | 77.4% | 44.4% |
+| - | Qwen2-7B | 79.9% | 44.2% |
+| - | Qwen2-Math-7B | 80.4% | 50.4% |
+| SIaM Yu et al. (2024) | Qwen-2-Math-Base | 81.5% | 50.0% |
+| - | Internlm2-math-plus-7B | 84.0% | 54.4% |
+| OMI2 Li et al. (2025) | Qwen2.5-Coder-7B | 84.1% | 72.3% |
+| CODEI/O++ Li et al. (2025) | Qwen2.5-Coder-7B | 85.7% | 72.1% |
+| PyEdu Li et al. (2025) | Qwen2.5-Coder-7B | 85.8% | 71.4% |
+| CODEI/O Li et al. (2025) | Qwen2.5-Coder-7B | 86.4% | 71.9% |
+| OC-SFT-1 Li et al. (2025) | Qwen2.5-Coder-7B | 86.7% | 70.9% |
+| WI Li et al. (2025) | Qwen2.5-Coder-7B | 87.0% | 71.4% |
+| WI (Full) Li et al. (2025) | Qwen2.5-Coder-7B | 87.0% | 71.1% |
+| OMI2 (Full) Li et al. (2025) | Qwen2.5-Coder-7B | 88.5% | 73.2% |
+| - | DeepSeekMath-7B-RL | 88.2% | 51.7% |
+| CoMAT Leang et al. (2024) | GPT-4 | 93.7% | - |
+| CoT Ranaldi et al. (2025) | GPT-4 | 94.5% | - |
+| FCoT Lyu et al. (2023) | GPT-4 | 95.0% | - |
+| MathPrompter Imani et al. (2023) | GPT-4 | 95.6% | - |
+| QuaSAR Radford et al. (2018) | GPT-4 | 96.5% | - |
+| MathDivide Srivastava and Gandhi (2024) | GPT-4 | 96.8% | - |
+| SoT (Ours) | GPT-4o-mini | 96.0% | 79.1% |
+Experimental Results: As shown in Table 2, SoT achieved 96.0% on GSM8K and 79.1% on MATH on GPT-4o-mini.
+On GSM8K, SoT’s performance is close to the best result achieved by GPT-4 (e.g., MathDivide’s 96.8%) Chen et al.
+(2025) and outperforms the best 7B model (e.g., OMI2 Full’s 88.5%) Chen et al.
+(2025).
+On MATH, SoT’s 79.1% significantly outperforms the best mainstream 7B models (e.g., OMI2 Full’s 73.2%).
+This result indicates that SoT can achieve reasoning capabilities close to those of closed-source large models on lightweight models, significantly narrowing the performance gap between open-source models and GPT-4, validating its superiority in complex mathematical reasoning.
+A.4 Path Counting and Token-Cost Metrics
+To align the notion of “paths” across different methods, we adopt a unified definition.
+For a method and a sample , let denote the number of candidate reasoning trajectories generated on that contain a complete chain of reasoning and yield a directly scorable answer.
+The horizontal axis in our plots reports
+which reflects how many full candidate solutions, on average, the method explores per problem.
+Under this formulation:
+CoT: Each problem produces exactly one chain-of-thought, hence .
+CoT-SC(): Each problem samples independent CoT trajectories followed by majority voting, thus .
+GoT(): During multi-round graph search, each invocation that produces a complete candidate solution is counted as a path.
+Internal graph-expansion calls that generate intermediate nodes without a final answer do not count as paths, though their token usage is included in the bubble size.
+AoT(): Along the Markov chain, solving each of produces a directly scorable candidate answer.
+Hence each solve() call is counted as one path, and approximates the average number of solver calls per Markov chain.
+SoT(): In SoT, LLMRESOLVE generates syzygies , each constituting a full structured reasoning chain capable of producing an answer.
+Therefore the number of candidate paths is exactly the number of syzygies:
+The auxiliary calls used during the freeness stage do not count as paths, but all of their token usage is included in the overall cost.
+The bubble area in the figure corresponds to the average tokens per problem, computed as the total number of tokens consumed by all LLM calls (generation, intermediate analysis, candidate solutions, and scoring), averaged over the evaluation set.
+This metric reflects the overall reasoning cost of each method.